OpenStudy (anonymous):

the line through (6,-4) with slope 2/3 in point-slope. show work please, thank you the line through (6,-4) with slope 2/3 in point-slope. show work please, thank you @Mathematics

OpenStudy (kainui):

y=mx+b m=slope = 2/3 we can solve for b by plugging in that point to the formula with the slope 6=(2/3)(-4)+b b=26/3 so you know your equation is y=2/3x+26/3